please help, it my homework.
log5to power 125 + log4(1/16)+log2 to power 32 -log8 to power 8.

Answer by lwsshak3(11628)

log5to power 125 + log4(1/16)+log2 to power 32 -log8 to power 8
**
log5(125)=log5(5^3)=3 (base raised to a power=power)
log4(1/16)=log4(1/4^2)=log4(4^-2)=-2
log2 to power 32=log2(32)=log2(2^5)=5
log8 to power 8=log8(8)=1 (log of base=base)
..
log5to power 125 + log4(1/16)+log2 to power 32 -log8 to power 8
=3-2+5+1=7